What is the energy of the electron in the n=1 level on an atom with an atomic number of 39?
1
Expert's answer
2017-09-21T03:31:07-0400


E(n=1) = - (m*e^4*Z^2)/(8e0^2*h^2*n^2) = - (9.1*10^-31kg*(1.6*10^-19C)^4*39^2)/(8*(8.85*10^-12F/m)^2*(6.62*10^-34Jsec)^2*1^2) = 3.3*10^-15 J